Find the vector joining the points P(2, 3, 0) and Q(-1, -2, -4) directed from P to Q.



Select one:



A. 3i-5j-4k




B. 3i-5j+4k




C. 3i+5j-4k




D. -3i-5j-4k


1
Expert's answer
2022-04-04T15:43:04-0400

"P=\left( 2,3,0 \right) ,Q=\left( -1,-2,-4 \right) \\PQ=\left( -1-2,-2-3,-4-0 \right) =\left( -3,-5,-4 \right) =-3i-5j-4k\\D\,\,is\,\,correct"